On October 29, Lobo Co. began operations by purchasing razors for resale. The razors have a 90-day warranty. When a razor is returned, the company discards it and mails a new one from the Merchandise Inventory to the customer. The company's cost per new razor is $14 and its retail selling price is $70. The company expects warranty costs to equal 5% of dollar sales. The following transactions occurred.

Nov. 11 Sold 70 razors for $4,900 cash.
30 Recognized warranty expense related to November sales with an adjusting entry.
Dec. 9 Replaced 14 razors that were returned under the warranty.
16 Sold 210 razors for $14,700 cash.
29 Replaced 28 razors that were returned under the warranty.
31 Recognized warranty expense related to December sales with an adjusting entry.
Jan. 5 Sold 140 razors for $9,800 cash.
17 Replaced 33 razors that were returned under the warranty.
31 Recognized warranty expense related to January sales with an adjusting entry.

1. Prepare journal entries to record above transactions and adjustments.

Options to choose from:

  • Accounts payable
  • Accrued payroll payable
  • Bonus payable
  • Cash
  • Cost of goods sold
  • Deferred income tax liability
  • Earned services revenue
  • Earned ticket revenue
  • Employee benefits plan payable
  • Employee bonus expense
  • Employee fed. inc. taxes payable
  • Employee life insurance payable
  • Employee medical insurance payable
  • Employee union dues payable
  • Estimated warranty liability
  • Federal unemployment taxes payable
  • FICA—Medicare taxes payable
  • FICA—Social sec. taxes payable
  • Income taxes expense
  • Income taxes payable
  • Interest expense
  • Interest payable
  • Merchandise inventory
  • Notes payable
  • Payroll taxes expense
  • Salaries expense
  • Salaries payable
  • Sales
  • Sales salaries expense
  • Sales taxes payable
  • State unemployment taxes payable
  • Unearned services revenue
  • Unearned ticket revenue
  • Vacation benefits expense
  • Vacation benefits payable
  • Wages expense
  • Warranty expense
Answer #1
1.1
Date General Journal Debit Credit
11-Nov Cash 4,900
Sales 4,900
11-Nov Cost of goods sold 980 =70*14
Merchandise inventory 980
30-Nov Warranty expense 245 =4900*5%
Estimated warranty liability 245
9-Dec Estimated warranty liability 196 =14*14
Merchandise inventory 196
16-Dec Cash 14,700
Sales 14,700
16-Dec Cost of goods sold 2,940 =210*14
Merchandise inventory 2,940
29-Dec Estimated warranty liability 392 =28*14
Merchandise inventory 392
31-Dec Warranty expense 735 =14700*5%
Estimated warranty liability 735
1.2
Date General Journal Debit Credit
5-Jan Cash 9,800
Sales 9,800
5-Jan Cost of goods sold 1,960 =140*14
Merchandise inventory 1,960
17-Jan Estimated warranty liability 462 =33*14
Merchandise inventory 462
31-Jan Warranty expense 490 =9800*5%
Estimated warranty liability 490